# Mesa-Extrapolation: A Weave Position Encoding Method for Enhanced Extrapolation in LLMs

### v1

@@ 1 Introduction
Large Language Models (LLMs) with their powerful in-context learning capabilities Brown et al. (2020) offer versatile solutions to a wide-range of intelligent applications. However, one pressing challenge, the extrapolation problem Delétang et al. (2022) Zhang et al. (2022), dictates that the inference ability of LLMs sharply declines beyond their max training lengths, imposing a serious limitation on applications with long inputs. An naive solution is to extend the length of training samples. However, the inherent quadratic complexity of calculations presents practical challenges, demanding more resources, longer training time and higher cost.
Positional encoding (PE) has become a pivotal component of Transformer architecture, to compensate for the overlooking of position information by the attention mechanism. In the realm of extrapolation capability, PE is considered as a key factor influencing the extrapolating ability of LLMs. A few PE approaches, such as the popular RoPE Su et al. (2023) and ALiBi Press et al. (2021), claim to offer improved extrapolation capabilities and have gained widespread usage in industrial applications. Meanwhile, a counter-narrative has emerged. Some works demonstrate that Transformer can achieve better extrapolation capabilities by removing position encoding (NoPE) Kazemnejad et al. (2023), and contend that the mask already plays a significant role in capturing position information.
Inspired by Kazemnejad et al. (2023), we conduct a thorough investigation of the extrapolation problem by designing a specific Transformer model for this purpose and presenting detailed theoretical analysis. To the best of our knowledge, this is the first theoretical endeavor to understand the inner workings of extrapolation. We first elucidate the cause of NoPE’s failure when input exceeds the effective window length in Theorem 3.1, and analyse PE’s failure in 3.2. Building upon Theorem 3.2, we prove in Theorem 3.3 that through meticulous weave position, (i.e., weave PE), it is feasible to achieve extrapolation beyond the effective window length.
@@ 2 Background
Since the self-attention mechanism itself does not contain position information, PE components have become an integral part of the Transformer architecture. Cmmon choices for PE are either absolute, where each absolute position (e.g., $1,2,3,...$) is directly represented, or relative, where the distance between tokens is used as positional information. Absolute Position Embedding (APE) Vaswani et al. (2017) embeds each absolute position $i$ into position vector $\bm{p}_{i}$ and adds word embeddings to their corresponding $\bm{p}_{i}$, before feeding them to the LLMs. However, the extrapolation ability is limited because it cannot generalize to unseen positions. RoPE Su et al. (2023) rotates the query and key vectors with an angle proportional to their absolute positions, so the attention dot production only depends on relative distance between tokens, providing a relative positional encoding. T5’s Relative Bias first maps the relative distance ($i-j$) between tokens at positions $i$ and $j$ to a scalar bias value $b=f(i-j)$, where $f$ is a lookup table. ALiBi is similar to T5’s Relative Bias but instead subtracts a scalar bias from the attention score Press et al. (2021)(refer to Appendix D.2 for more details). Recent works such as ReRoPE and Leaky-ReRoPE Su (2023b) achieve effective extrapolation without fine-tuning by meticulously weaving relative positions of RoPE. We refer to this class of methods that achieve extrapolation by weaving the relative positions of PE without fine-tuning as Weave PE.
Kazemnejad et al. (2023) indicates that the decoder-only transformer with position encoding removed (NoPE) demonstrates stronger extrapolation capabilities. Furthermore, it theoretically shows that a specific transformer model can get relative and absolute positional information, even in the absence of PE. Haviv et al. (2022) also demonstrates NoPE achieves comparative performance to standard Transformer models. These new studies pose a key challenge regarding the choice of whether using PE or not in Transformer architecture (refer to Appendix A for more related works).
@@ 3.1 Problem Definition
$\small\langle\bm{q}_{t},\bm{k}_{i}\rangle:=\bm{q}_{t}^{T}\bm{k}_{i}$
@@
For example, ReRoPE Su (2023b) can be considered as an example of weave PE, with its $\mathcal{W}$ function defined as follows:
@@
$\small\langle\bm{q}_{t},\bm{k}_{i}\rangle:=f_{\mathrm{RoPE}}(\bm{q}_{t},\bm{k}_{i},t-i))=\bm{q}_{t}^{T}R^{(t-i)\theta}\bm{k}_{i}$
@@ 3.2 Motivation
Chen et al. (2023) explores the evolution of hidden state values and reveals a noticable phenomenon: as the position increases, the hidden state values will explode. This finding appears consistent with observed failures in extrapolation. Through probe experiments (refer to Appendix F), we investigate the alterations in hidden state values across various positions and layers. Results show a significant shift in the hidden state’s value range upon surpassing the effective window. Interestingly, when employing extrapolation techniques, hidden state values exhibit noticeable suppression. This indicates that the effective range of hidden state values likely lies within a specific threshold. When the position exceeds the effective window length, the hidden state values surpass this threshold, resulting in extrapolation failures.
Previous work Kazemnejad et al. (2023) utilizes a constructive approach. By constructing the Transformer’s weights, it enables the first and second layers to independently generate position information. Drawing inspiration from this, we endeavor to construct a Transformer model capable of mirroring this observation.
@@ 3.3 Theoretical Analysis
Based on our observations, thresholds in different dimensions can serve as either upper or lower bounds. For simplicity, we treat thresholds solely as lower bounds and make the following assumptions:
@@
Building upon this assumption, theoretical results for NoPE exceeding the effective window length are as follows:
@@
Through Theorem 3.1 and Theorem 3.2, we formulate pertinent theoretical models for NoPE and PE, respectively, shedding light on the intricate relationship between extrapolation and the effective window length. Building upon these findings, Theorem 3.3 delves deeper into the realm of explicit PE, revealing that a well-designed weave PE scheme can effectively broaden the original effective window length. The theorems show the existence of an effective length $M$, which is typically related to the maximum training window length of the LLM. Furthermore, within the proof of Theorem 3.3, our results indicate that $N\ll M$, consistent to experimental findings in Su (2023b), where although the extrapolation position can theoretically start from 2048, the best extrapolation starting position is 512. More experimental parameters also validate this setting (refer to Appendix B.2).
@@ 3.4 Validating Extrapolation Using Observed Thresholds
Further, we design probe experiments (refer to Appendix F.3 for more results) to validate the observed phenomena and our theorems, as shown in Figure 2. From Figure 2, two observations are noted: Firstly, for hidden state values of the same dimension, the first layer undergoes minimal change, while the second layer exhibits a more pronounced transition. Exceeding the threshold implies extrapolation failure. This observation aligns with our theoretical model construction, where the first layer primarily refines positional information, with significant signal changes occurring in the second layer, as demonstrated in Theorem 3.2. Secondly, based on observational thresholds, when the length of the input sequence is around 12k, the values of hidden state corresponding to Dynamic-NTK Liu et al. (2023) surpass the threshold, implying extrapolation failure. Conversely, for ReRoPE Su (2023b), extrapolation succeeds. These two predictive outcomes corroborate with subsequent experimental results.
@@ 4.1 Stair PE
$\small\langle\bm{q}_{t},\bm{k}_{i}\rangle:=f_{\mathrm{StairRoPE}}(\bm{q}_{t},\bm{k}_{i},t-i)=f_{\mathrm{RoPE}}(\bm{q}_{t},\bm{k}_{i},\mathcal{W}(t-i))=\bm{q}_{t}^{T}R^{\mathcal{W}(t-i)\theta}\bm{k}_{i}$
Taken a sequence of length $10$ as an example, the right subplot of Figure 1 shows that the relative positions generated by PE (Both RoPE and ALiBi) are linear, while those generated by Stair PE (here $N=4$ and $E=2$) are non-linear. Since weave PE changes the linear relative position, it has to calculate the attention matrix more than once Su (2023b). Compared with ReRoPE (detailed on Appendix B.3), Stair PE provides a finer-grained extrapolated positions. Meanwhile, compared with Leaky-ReRoPE, Stair PE reuses known positions, reducing possible generalization errors. We also provide an ablation experiment to compare these weave PE methods (refer to Appendix C.6).
While our work was conducted independently, we note that Jin et al. (2024) have recently explored a similar idea through flooring the original positions and obtaining the relative position matrix with grouped self-attention. Although the two parallel thought processes are different, under certain conditions their formulations are equivalent (refer to Appendix G). Consequently, Self-Extend can be categorized as a Weave PE method. Our proposed Chunk-based Triangular Attention Matrix (detailed in Section 4.2) and its corresponding theoretical properties (Section 4.4) are also applicable to this parallel approach.
@@ 5 Experiments
In this section, we validate the performance of Mesa-Extrapolation through experiments measured over multiple metrics. We choose GovReport Huang et al. (2021), Pile Gao et al. (2020), LongBench Bai et al. (2023), and LongEval Krishna et al. (2023) datasets, and also generate a passkey dataset, which has been integrated in the code warehouse. More experimental details and results are referred to Appendix B and C.
Since our method is completely free plug-in and does not require fine-tuning, we choose methods of this type for comparison, including: model self (Origin), ReRoPE Su (2023b), Leaky-ReRoPE Su (2023b), Dynamic-NTK Liu et al. (2023), LM-Infinite Han et al. (2023), and Streaming-LLM Xiao et al. (2023).
We evaluate Mesa-Extrapolation using three prominent LLM families: LLaMA Touvron et al. (2023a) Touvron et al. (2023b) (including LLaMA-3B (Open-LLaMA-3B), LLaMA2-7B-Chat, and Vicuna-13B-V1.3), MPT Team (2023) (including MPT-7B), and PyThia Biderman et al. (2023) (including PyThia-6.9B and PyThia-12B). Notably, LLaMA and PyThia incorporate RoPE Su et al. (2023), whereas MPT employs ALiBi Press et al. (2021) – two of the most influential PE techniques in recent research. Furthermore, we validated our approach using the Phi-3-mini-128k-instruct model Microsoft (2024) (refer to Appendix C.9).
@@ 5.3 Evaluation on Summary of Tasks
We conduct a summary task using the GovReport dataset and employ ROUGE ROUGE (2004) (ROUGE-1/2/L) as evaluation metrics. ROUGE assess overlapping N-grams by comparing the generated text with reference answers.
@@ 6 Conclusion
Limitations. Mesa-Extrapolation is a plug-and-play method that does not require additional fine-tuning. However, previous work, "NTK-aware" ntk (2023) also shows that applying further fine-tuning to plug-in extrapolation is possible. Therefore exploring fine-tuning based on Mesa-Extrapolation can be an interesting next step.
@@ Appendix A Related Work
Extrapolation for PE A distinct line of research is dedicated to refining PE for enhanced extrapolation capabilities. Notable contributions include the introduction of Rotary Position Encoding (RoPE) Su et al. (2023), implementing relative PE through absolute position information. Similarly, Press et al. (2021) proposes a novel PE method ALiBi, fusing position information by directly introducing the relative position distance term in dot multiplication. In addition, there are other PE methods such as absolute position embedding (APE) Vaswani et al. (2017) and T5’s Relative PE Raffel et al. (2020).
@@
Another school of research focuses on enhancing the extrapolation performance of existing LLMs, categorized by whether to include further training. The first subcategory involves further fine-tuning, enlarging the effective window length by training LLMs on longer input texts. Chen et al. (2023) demonstrates that Position Interpolation (PI) method has a superior fine-tune effect, resulting in extended extrapolation capabilities with fewer fine-tuning steps. Mohtashami & Jaggi (2023) utilizes a new landmark token to represent individual input blocks, and enables model to select relevant blocks by further training, allowing for longer extrapolation. Zhang et al. (2024) follows similar idea by training a special token. Bertsch et al. (2023) proposes Unlimiformer, a k-nearest-neighbor (kNN) indexed encoder-decoder Transformer, enhancing efficiency by retrieving top-k keys for each decoder layer. Despite claiming support for decoder-only Transformer, differences in retrieval results across decoder layers may lead to potential failure. Tworkowski et al. (2023) introduces Focused Transformer (FOT), utilizing a contrastive learning-inspired training process to enhance the (key, value) space structure and enable effective context extension. The second subcategory explores methods that require no fine-tuning yet offer improved extrapolation through plug-ins. Su (2023b) introduces Rectified Rotary Position Embeddings (ReRoPE), a method that rectifies extrapolated relative positions based on RoPE within a specified interval to extend the effective window length. In addition, Leaky-ReRoPE offers an alternative by allowing a controlled leakage of position extrapolation within an interval. Both approaches are well-suited for LLaMA model families Touvron et al. (2023b) Touvron et al. (2023a), eliminating the need for fine-tuning. These methods do suffer from some drawbacks, such as twofold increase in memory consumption and longer inference time. We refer to this class of methods that achieve extrapolation by weaving the relative positions of PE without fine-tuning as Weave PE. We also notice that, recent work, InfLLM Xiao et al. (2024), proposes an additional memory units, which lookup token-relevant units for attention computation. In addition, it uses a modified encoding scheme similar to ReRoPE to achieve longer extrapolation. It is worth noting that the methods of the weave PE class can be applied seamlessly to these new designs. However, there is still no theory to explain why the methods of weave PE class can work. In a different line of inquiry, ntk (2023) proposes the "NTK-aware" method by drawing from the Neural Tangents (NTK) idea, which explores the high-frequency extrapolation and low-frequency interpolation concept for RoPE. Based on "NTK-aware", recent works bloc97 (2023), Peng et al. (2023), Roziere et al. (2023), Liu et al. (2023) perform further fine-tuning for optimal results.
@@
Han et al. (2023) proposes LM-Infinite, which employs a $\Lambda$-shaped mask to prevent surpassing the effective window length by discarding central tokens. However, this design choice inevitably results in a loss of information. Likewise, Streaming-LLM Xiao et al. (2023) adopts a similar strategy to avoid surpassing the effective window length by discarding a portion of input tokens.
@@
Extrapolation for NoPE There also exists a counter perspective asserting that the position information of an input sequence can be perceived without utilizing PE. In a comprehensive experimental comparison presented in Kazemnejad et al. (2023), the decoder-only Transformer is shown to exhibit superior extrapolation properties with no position encoding (NoPE). Haviv et al. (2022) conjectures that causal attention enables the Transformer to infer position information without the help of PE, demonstrating its comparable performance with standard Transformer models through probing experiments. These new studies pose a key challenge regarding the choice of whether using PE or not in Transformer architecture.
@@
Theorems for Extrapolation We mainly focus on Transformer architecture. Although the Transformer architecture is considered as a "black box", there are ongoing efforts trying to explain its inner workings. Von Oswald et al. (2023) begins by providing a specific weight construction that elucidates the mechanistic understanding of in-context learning within optimized Transformers. Lindner et al. (2023) introduces interpretability for Transformers through programming, allowing the derivation of a program Transformer architecture tailored for specific tasks. In Han et al. (2023), the authors conclude that long-distance attention logits will explode. However, this does not clarify the relationship between the effective window length and the existing supremum. It merely indicates an increase in the supremum, without directly implying an increase in the obtained actual attention logits. In Kazemnejad et al. (2023), inspired by programming Transformers Lindner et al. (2023), the authors construct a specific Transformer with predefined matrix parameters, theoretically demonstrating that NoPE can recover both absolute and relative position information even without the use of PE. However, these results fall short in explaining the underlying reasons for the failure of NoPE extrapolation. Built upon the entropy increase theory, Han et al. (2023) proves that with the expansion of input length, the entropy of attention experiences a corresponding increase. Additionally, Su (2021) introduces a scaling factor, denoted as $log_{\mathrm{train{-}len}}(n)$, which serves to mitigate entropy increase. Nevertheless, their theorems do not show that extrapolation fails beyond the effective window length.

### v3

@@ 1 Introduction
Large Language Models (LLMs) with their powerful in-context learning capabilities [8] offer versatile solutions to a wide-range of intelligent applications. However, one pressing challenge, the extrapolation problem [10] [48], dictates that the inference ability of LLMs sharply declines beyond their max training lengths, imposing a serious limitation on applications with long inputs. An naive solution is to extend the length of training samples. However, the inherent quadratic complexity of calculations presents practical challenges, demanding more resources, longer training time and higher cost.
Positional encoding (PE) has become a pivotal component of Transformer architecture, to compensate for the overlooking of position information by the attention mechanism. In the realm of extrapolation capability, PE is considered as a key factor influencing the extrapolating ability of LLMs. A few PE approaches, such as the popular RoPE [36] and ALiBi [29], claim to offer improved extrapolation capabilities and have gained widespread usage in industrial applications. Meanwhile, a counter-narrative has emerged. Some works demonstrate that Transformer can achieve better extrapolation capabilities by removing position encoding (NoPE) [20], and contend that the mask already plays a significant role in capturing position information.
Inspired by [20], we conduct a thorough investigation of the extrapolation problem by designing a specific Transformer model for this purpose and presenting detailed theoretical analysis. To the best of our knowledge, this is the first theoretical endeavor to understand the inner workings of extrapolation. We first elucidate the cause of NoPE’s failure when input exceeds the effective window length in Theorem 3.1, and analyse PE’s failure in 3.2. Building upon Theorem 3.2, we prove in Theorem 3.3 that through meticulous weave position, (i.e., weave PE), it is feasible to achieve extrapolation beyond the effective window length.
@@ 2 Background
Since the self-attention mechanism itself does not contain position information, PE components have become an integral part of the Transformer architecture. Cmmon choices for PE are either absolute, where each absolute position (e.g., $1,2,3,...$) is directly represented, or relative, where the distance between tokens is used as positional information. Absolute Position Embedding (APE) [41] embeds each absolute position $i$ into position vector $\bm{p}_{i}$ and adds word embeddings to their corresponding $\bm{p}_{i}$, before feeding them to the LLMs. However, the extrapolation ability is limited because it cannot generalize to unseen positions. RoPE [36] rotates the query and key vectors with an angle proportional to their absolute positions, so the attention dot production only depends on relative distance between tokens, providing a relative positional encoding. T5’s Relative Bias first maps the relative distance ($i-j$) between tokens at positions $i$ and $j$ to a scalar bias value $b=f(i-j)$, where $f$ is a lookup table. ALiBi is similar to T5’s Relative Bias but instead subtracts a scalar bias from the attention score [29](refer to Appendix D.2 for more details). Recent works such as ReRoPE and Leaky-ReRoPE [35] achieve effective extrapolation without fine-tuning by meticulously weaving relative positions of RoPE. We refer to this class of methods that achieve extrapolation by weaving the relative positions of PE without fine-tuning as Weave PE.
[20] indicates that the decoder-only transformer with position encoding removed (NoPE) demonstrates stronger extrapolation capabilities. Furthermore, it theoretically shows that a specific transformer model can get relative and absolute positional information, even in the absence of PE. [15] also demonstrates NoPE achieves comparative performance to standard Transformer models. These new studies pose a key challenge regarding the choice of whether using PE or not in Transformer architecture (refer to Appendix A for more related works).
@@ 3.1 Problem Definition
$\small\langle\bm{q}_{t},\bm{k}_{i}\rangle:=\bm{q}_{t}^{T}\bm{k}_{i}.$
@@
For example, ReRoPE [35] can be considered as an example of weave PE, with its $\mathcal{W}$ function defined as follows:
@@
$\small\langle\bm{q}_{t},\bm{k}_{i}\rangle:=f_{\mathrm{RoPE}}(\bm{q}_{t},\bm{k}_{i},t-i))=\bm{q}_{t}^{T}R^{(t-i)\theta}\bm{k}_{i}.$
@@ 3.2 Motivation
[9] explores the evolution of hidden state values and reveals a noticable phenomenon: as the position increases, the hidden state values will explode. This finding appears consistent with observed failures in extrapolation. Through probe experiments (refer to Appendix F), we investigate the alterations in hidden state values across various positions and layers. Results show a significant shift in the hidden state’s value range upon surpassing the effective window. Interestingly, when employing extrapolation techniques, hidden state values exhibit noticeable suppression. This indicates that the effective range of hidden state values likely lies within a specific threshold. When the position exceeds the effective window length, the hidden state values surpass this threshold, resulting in extrapolation failures.
Previous work [20] utilizes a constructive approach. By constructing the Transformer’s weights, it enables the first and second layers to independently generate position information. Drawing inspiration from this, we endeavor to construct a Transformer model capable of mirroring this observation.
@@ 3.3 Theoretical Analysis
In a multi-layer neural network, each layer’s outputs, a.k.a hidden state values $o$, become the inputs for the subsequent layer. To maintain stable network behavior, these values must remain within a reasonable range. We define this observable boundary as the threshold $\mathcal{H}$. This threshold can be either an upper bound or a lower bound. For our analysis, we focus on the lower bound of this threshold. A successful extrapolation occurs when a large model consistently generates accurate next tokens for a long input sequence. Conversely, a failed extrapolation happens when the model produces incorrect or nonsensical next tokens. Based on these definitions, we make the following assumptions:
@@
These assumptions indicate that by observing whether the hidden state value $o$ in this dimension exceed the threshold $\mathcal{H}$, we can predict whether the large model’s extrapolation has failed. Building upon these assumptions, theoretical results for NoPE exceeding the effective window length are as follows:
@@
Through Theorem 3.1 and Theorem 3.2, we formulate pertinent theoretical models for NoPE and PE, respectively, shedding light on the intricate relationship between extrapolation and the effective window length. Building upon these findings, Theorem 3.3 delves deeper into the realm of explicit PE, revealing that a well-designed weave PE scheme can effectively broaden the original effective window length. The theorems show the existence of an effective length $M$, which is typically related to the maximum training window length of the LLM. Furthermore, within the proof of Theorem 3.3, our results indicate that $N\ll M$, consistent to experimental findings in [35], where although the extrapolation position can theoretically start from 2048, the best extrapolation starting position is 512. More experimental parameters also validate this setting (refer to Appendix B.2).
@@ 3.4 Validating Extrapolation Using Observed Thresholds
Further, we design probe experiments (refer to Appendix F.3 for more results) to validate the observed phenomena and our theorems, as shown in Figure 2. From Figure 2, two observations are noted: Firstly, for hidden state values of the same dimension, the first layer undergoes minimal change, while the second layer exhibits a more pronounced transition. Exceeding the threshold implies extrapolation failure. This observation aligns with our theoretical model construction, where the first layer primarily refines positional information, with significant signal changes occurring in the second layer, as demonstrated in Theorem 3.2. Secondly, based on observational thresholds, when the length of the input sequence is around 12k, the values of hidden state corresponding to Dynamic-NTK [24] surpass the threshold, implying extrapolation failure. Conversely, for ReRoPE [35], extrapolation succeeds. These two predictive outcomes corroborate with subsequent experimental results.
@@ 4.1 Stair PE
$\small\langle\bm{q}_{t},\bm{k}_{i}\rangle:=f_{\mathrm{StairRoPE}}(\bm{q}_{t},\bm{k}_{i},t-i)=f_{\mathrm{RoPE}}(\bm{q}_{t},\bm{k}_{i},\mathcal{W}(t-i))=\bm{q}_{t}^{T}R^{\mathcal{W}(t-i)\theta}\bm{k}_{i}.$
Taken a sequence of length $10$ as an example, the right subplot of Figure 1 shows that the relative positions generated by PE (Both RoPE and ALiBi) are linear, while those generated by Stair PE (here $N=4$ and $E=2$) are non-linear. Since weave PE changes the linear relative position, it has to calculate the attention matrix more than once [35]. Compared with ReRoPE (detailed on Appendix B.3), Stair PE provides a finer-grained extrapolated positions. Meanwhile, compared with Leaky-ReRoPE, Stair PE reuses known positions, reducing possible generalization errors. We also provide an ablation experiment to compare these weave PE methods (refer to Appendix C.6).
While our work was conducted independently, we note that [19] have recently explored a similar idea through flooring the original positions and obtaining the relative position matrix with grouped self-attention. Although the two parallel thought processes are different, under certain conditions their formulations are equivalent (refer to Appendix G). Consequently, Self-Extend can be categorized as a Weave PE method. Our proposed Chunk-based Triangular Attention Matrix (detailed in Section 4.2) and its corresponding theoretical properties (Section 4.4) are also applicable to this parallel approach.
@@ 5 Experiments
In this section, we validate the performance of Mesa-Extrapolation through experiments measured over multiple metrics. We choose GovReport [18], Pile [12], LongBench [3], and LongEval [21] datasets, and also generate a passkey dataset, which has been integrated in the code warehouse. More experimental details and results are referred to Appendix B and C.
Since our method is completely free plug-in and does not require fine-tuning, we choose methods of this type for comparison, including: model self (Origin), ReRoPE [35], Leaky-ReRoPE [35], Dynamic-NTK [24], LM-Infinite [14], and Streaming-LLM [45].
We evaluate Mesa-Extrapolation using three prominent LLM families: LLaMA [38] [39] (including LLaMA-3B (Open-LLaMA-3B), LLaMA2-7B-Chat, and Vicuna-13B-V1.3), MPT [37] (including MPT-7B), and PyThia [5] (including PyThia-6.9B and PyThia-12B). Notably, LLaMA and PyThia incorporate RoPE [36], whereas MPT employs ALiBi [29] – two of the most influential PE techniques in recent research. Furthermore, we validated our approach using the Phi-3-mini-128k-instruct model [25] (refer to Appendix C.9).
@@ 5.3 Evaluation on Summary of Tasks
We conduct a summary task using the GovReport dataset and employ ROUGE [31] (ROUGE-1/2/L) as evaluation metrics. ROUGE assess overlapping N-grams by comparing the generated text with reference answers.
@@ 6 Conclusion
Limitations. Mesa-Extrapolation is a plug-and-play method that does not require additional fine-tuning. However, previous work, "NTK-aware" [1] also shows that applying further fine-tuning to plug-in extrapolation is possible. Therefore exploring fine-tuning based on Mesa-Extrapolation can be an interesting next step. Due to limitations of resources, we have not yet validated our method at longer lengths.
@@ Appendix A Related Work
Extrapolation for PE A distinct line of research is dedicated to refining PE for enhanced extrapolation capabilities. Notable contributions include the introduction of Rotary Position Encoding (RoPE) [36], implementing relative PE through absolute position information. Similarly, [29] proposes a novel PE method ALiBi, fusing position information by directly introducing the relative position distance term in dot multiplication. In addition, there are other PE methods such as absolute position embedding (APE) [41] and T5’s Relative PE [30].
@@
Another school of research focuses on enhancing the extrapolation performance of existing LLMs, categorized by whether to include further training. The first subcategory involves further fine-tuning, enlarging the effective window length by training LLMs on longer input texts. [9] demonstrates that Position Interpolation (PI) method has a superior fine-tune effect, resulting in extended extrapolation capabilities with fewer fine-tuning steps. [26] utilizes a new landmark token to represent individual input blocks, and enables model to select relevant blocks by further training, allowing for longer extrapolation. [47] follows similar idea by training a special token. [4] proposes Unlimiformer, a k-nearest-neighbor (kNN) indexed encoder-decoder Transformer, enhancing efficiency by retrieving top-k keys for each decoder layer. Despite claiming support for decoder-only Transformer, differences in retrieval results across decoder layers may lead to potential failure. [40] introduces Focused Transformer (FOT), utilizing a contrastive learning-inspired training process to enhance the (key, value) space structure and enable effective context extension. The second subcategory explores methods that require no fine-tuning yet offer improved extrapolation through plug-ins. [35] introduces Rectified Rotary Position Embeddings (ReRoPE), a method that rectifies extrapolated relative positions based on RoPE within a specified interval to extend the effective window length. In addition, Leaky-ReRoPE offers an alternative by allowing a controlled leakage of position extrapolation within an interval. Both approaches are well-suited for LLaMA model families [39] [38], eliminating the need for fine-tuning. These methods do suffer from some drawbacks, such as twofold increase in memory consumption and longer inference time. We refer to this class of methods that achieve extrapolation by weaving the relative positions of PE without fine-tuning as Weave PE. We also notice that, recent work, InfLLM [44], proposes an additional memory units, which lookup token-relevant units for attention computation. In addition, it uses a modified encoding scheme similar to ReRoPE to achieve longer extrapolation. It is worth noting that the methods of the weave PE class can be applied seamlessly to these new designs. However, there is still no theory to explain why the methods of weave PE class can work. In a different line of inquiry, [1] proposes the "NTK-aware" method by drawing from the Neural Tangents (NTK) idea, which explores the high-frequency extrapolation and low-frequency interpolation concept for RoPE. Based on "NTK-aware", recent works [7], [28], [32], [24] perform further fine-tuning for optimal results.
@@
[14] proposes LM-Infinite, which employs a $\Lambda$-shaped mask to prevent surpassing the effective window length by discarding central tokens. However, this design choice inevitably results in a loss of information. Likewise, Streaming-LLM [45] adopts a similar strategy to avoid surpassing the effective window length by discarding a portion of input tokens.
@@
Extrapolation for NoPE There also exists a counter perspective asserting that the position information of an input sequence can be perceived without utilizing PE. In a comprehensive experimental comparison presented in [20], the decoder-only Transformer is shown to exhibit superior extrapolation properties with no position encoding (NoPE). [15] conjectures that causal attention enables the Transformer to infer position information without the help of PE, demonstrating its comparable performance with standard Transformer models through probing experiments. These new studies pose a key challenge regarding the choice of whether using PE or not in Transformer architecture.
@@
Theorems for Extrapolation We mainly focus on Transformer architecture. Although the Transformer architecture is considered as a "black box", there are ongoing efforts trying to explain its inner workings. [42] begins by providing a specific weight construction that elucidates the mechanistic understanding of in-context learning within optimized Transformers. [23] introduces interpretability for Transformers through programming, allowing the derivation of a program Transformer architecture tailored for specific tasks. In [14], the authors conclude that long-distance attention logits will explode. However, this does not clarify the relationship between the effective window length and the existing supremum. It merely indicates an increase in the supremum, without directly implying an increase in the obtained actual attention logits. In [20], inspired by programming Transformers [23], the authors construct a specific Transformer with predefined matrix parameters, theoretically demonstrating that NoPE can recover both absolute and relative position information even without the use of PE. However, these results fall short in explaining the underlying reasons for the failure of NoPE extrapolation. Built upon the entropy increase theory, [14] proves that with the expansion of input length, the entropy of attention experiences a corresponding increase. Additionally, [33] introduces a scaling factor, denoted as $log_{\mathrm{train{-}len}}(n)$, which serves to mitigate entropy increase. Nevertheless, their theorems do not show that extrapolation fails beyond the effective window length.